Paranal Previous News
This page contains Paranal Telescope and Instrument news items for the previous years 2004 - 2005.
- September 19 to October 8, 2012: UT4/Yepun was out of operations due to a serious problem with the telescope hydraulics. The repair was completed on Oct. 7/8 and the telescope was recalibrated and handed over to operations in a very good shape.
- September 7, 2012: recoating of M1, M3 of UT3 completed
- June 11, 2012: recoating of M1, M3 of UT1 completed
- June 4, 2012: UT4 - Yepun: The UT4 Laser Guide Star (LGS) system will be unavailable for operations in January and February 2013. It will be replaced by a prototype laser for the Adaptive Optics Facility (AOF). This new laser system is expected to provide easier maintenance and operations, and the number of LGS nights available should significantly increase.
- March 21 - May 3: UT4 - Yepun: Because of preparatory work for the installation of the Adaptive Optics Facility, Yepun is not available for operations.
- November 17, 2011: recoating of M1, M3 of UT2 completed
- August 13, 2011: Start of VST Public Surveys with Early Science and Dry Runs in parallel to technical activities
- March 14 - May 4, 2011: Extended VISTA recovery period. A serious mechanical problem detected during the coating intervention requires an extended repair and preventive maintenance period. The VISTA telescope is expected to be back for on-sky tests by May 3/4, 2011.
- March 4 - April 7, 2011: Recoating of M1 and M2 of VISTA with aluminium.
- December, 2010:Commissioning of the VST started with all systems integrated. A smalltest CCD is installed in place of OMEGACAM (the science instrument).OMEGACAM will be installed at the VST in March 2011.
- November 23, 2010: recoating of M1 of UT1 completed
- September 28, 2010: recoating of M1 of UT3 completed
- November 3, 2009: recoating of M1 of UT2 completed
- September 8, 2009: recoating of M1 of UT4 completed
- June, 2009:
The last part of the VST, the M1 cell, arrived to Paranal this month. Engineering activities will now continuously ramp up throughout the second semester of 2009, aiming for a start of operations of the VST in 2010.
Photo G. Hüdepohl
- November 13, 2008: recoating of M1 of UT1 completed
- November, 2008: After a successful period of technical activities, the LGS facility is now fully offered with both NACO and SINFONI.
- August, 28, 2008: Following the LGS recovery actions which took place at the end of July, it has been decided to resume operations of SINFONI with the laser (service mode only). The use of the laser with NACO is still waiting for the results of a commissioning run scheduled for November 2008.Moreover, due to the intrinsic delays in getting the LGS system in operations, the commissioning of the seeing-enhancer mode with SINFONI has been postponed to early 2009. Updated news will be posted on this webpage when they become available.
- June, 26, 2008: Although substantial improvements were made to the LGS during the recent technical activities that took place at UT4, it was decided that an additional set of recovery actions was needed - to be scheduled end of July 2008. Until then, the LGSF remains unavailable for science operations.Please, continue to check this webpage for the latest news on the LGSF recovery activities.
- May, 23, 2008: LGSF recovery activities have been extended until mid-June 2008.As a result, the LGSF remains unavailable for scienceoperations. Depending on the success of the June LGS recovery,executions of some science programs in Service Mode couldresume by the end of July 2008.
News on the ongoing recovery activities will continue to be posted here.
- April, 30, 2008: The continued degradation of the performance, reliability, and availability of the Laser Guide Star Facility (LGSF) has forced to Observatory to take this facility out of operation as of April 1st, 2008. All service mode programs requiring the LGS are on hold. Visitor runs with the LGS will have to be cancelled.
After a deep analysis of the situation a detailed and comprehensive recovery plan has been developed by ESO together with the LGS partner institutes to remedy this situation. The plan was released on April 25 and foresees the implementation of immediate short-term actions for the period April - July to recover the LGSF performance as required for successful science operation (i.e., the performance as of June 2007).
A first critical milestone of the recovery plan is to have the LGSF operational for mid-May where a LGS commissioning together with NACO has been scheduled earlier.
In parallel to the implementation of the short-term plan a medium-term plan for the LGSF is developed to improve its reliability and performance on system level. This plan is expected to be released in October this year.
- April 1st, 2008: AMBER: Fringe tracking with FINITO is availabe to visitor mode PIs
- April 1st, 2008: HAWK-I: start of Science Operation
- July 2, 2007: recoating of M1 of UT1 successfully completed
- March 8, 2007: recoating of M2 of UT2 successfully completed
- January 2007: AT4 commissioned and accepted for operations
- December 5, 2006: UT2 recoating of M1/M3 completed
- September 5, 2006: UT4 recoating of M1/M3 completed
- May 12, 2006: First recoating of M2 of UT1 successfully completed
- December 16, 2005: UT1 recoating completed
- November 01, 2005: first light for AT3
- October 20, 2005: UT3 recoating completed
- June 22, 2005: UT2 recoating completed
- April 27, 2005: UT4 recoating completed
- February 2005: first fringes with two ATs with the VINCI and MIDI instruments
- February 2005: both VLTI Siderostats decommissioned
- January 2005: AT2 arrived on mountain top
- January 2004: AT1 arrived on mountain top
- December 21: VISIR: sensitivities in imaging and low resolution spectroscopy have been updated for phase 2 of P97. Also an M-band filter has been partially commissioned. See the VISIR news page.
- December 21: Start of a SINFONI upgrade: it includes replacement of the J, H, K and H+K filters, replceament of the pre-optics with better J-band transmission, replacement of mirrors with better surface quality ones to improve the line spread function in particular in the J-band. Re-commissioning is expected to take place on January 29 and 30.
- December 15: Publications of early science results of the XXL survey, which aimed at providing a well-defined sample of about 500 galaxy clusters out to a distance when the Universe was half of its current age: five new superclusters of galaxies, fewer distance clusters than expected, etc... The survey combines the XMM-Newton survey and with data obtained from ground based observatories, including from EFOSC2, FORS and MUSE.
- December 14: NACO is back in operation. All SM modes are operational, but some VM modes need some additional work. The camera wheel has recovered its functionality. Quadrant 4 has also been recovered; however, quadrant 2 (lower left) has not been recovered: 1 in 8 columns has no signal. See the NACO news.
- December 9: Observations with FORS and MUSE reveals the aftermath of a 360 million year collision between galaxies, including a peculiar young dwarf galaxy.
- December 2-17: A problem with an electronic board of SAXO impedes observations with SPHERE. See the SPHERE news.
- November 28: A 6.2 earthquake hit Paranal on November 27 at 18:00 local time: its epicenter was very close (within 30 km) to the observatory. After the necessary checks, most instruments were shown to work nominally afterwards. However, the UVES blue arm camera shows coma which reduces its spectral resolution.
- November 25: SPHERE obtained the most detailed image of the hypergiant star VY Canis Majoris.
- November 25: MUSE observations lead to several predictions regarding the appearance of the Refsdal supernova lensed by a galaxy cluster.
- November 23: First PIONIER service mode observations with the upgraded VLTI.
- November 20 - December 4: Maintenance of the VIRCAM cooling system and software upgrade.
- November 18: The UltraVISTA survey uncovered 574 new massive galaxies but none earlier than about obe billion years after the Big Bang.
- November 16: The Call for proposals for the VISIR Upgrade Science Verification has been issued. See the VISIR news page.
- November 11: Doppler tomography using several instruments, including UVES and X-shooter, over a twelve-years period details the shattered remains of an asteroid around a white dwarf.
- November 4-17: first, successful, commissioning run of GRAVITY.
- October 28: The VISTA Variables in the Via Lactea survey discovers a disc of young stars buried behind thick coulds in the Mily Way central Bulge.
- October 21: FLAMES finds hottest and most massive touching double star.
- October 7: Comparison between images obtained with SPHERE and previous NASA/ESA Hubble images revealed fast-moving wave-like features in the disc of the star AU Mic.
- October 1: Start of a major intervention on NACO to recover the camera wheel and improve the detector characteristics.
- September 4: The AG slide of X-shooter will have to be repaired in the coming months. See the X-shooter news.
- September 1: The first commissioning of GRAAL, part of the Adaptive Optics Facility, is completed. HAWK-I is back in operation.
- August 25: First fringes with the VLTI following the upgrade: MIDI as been removed from the laboratory, PIONIER now sits above FINITRO, the laboratory can now accomodate GRAVITY and MATISSE, new relay optic systems have been installed on the ATs, a new AT maintenance station has been commissioned, new hardware has been installed for GRAVITY, new delay line relay mirroes have been commissioned for new operation stations, the UT Coudé rooms have been prepared for the GRAVITY wavefron sensors (CIAO)
- August 14-26: Unstability of the temperature of both detectors lead to suspend observations with VISIR. See VISIR news page.
- August 12: Twilight flats obtained after the FORS2 LADC prism exchange have now been analysed, confirming that the structure clearly visible in the old data is gone. See the FORS2 News page.
- August 10: The Galaxy and Mass Assembly (GAMA) survey releases first data, obtained with VIRCAM and OmegaCAM, as well as NASA GALEX and WISE and ESA Herschel. It meausures the energy output of over 200,000 galaxies across as broad a wavelength range as possible.
- August 1: A Call for Letters of Intent for new public surveys with VISTA was released. See the VIRCAM News page.
- July 28: FLAMES: Spectral resolutions have improved following the update in the focussing procedure: see the FLAMES News page.
- July 19: MUSE: detector voltage configuration files for Channels 1 and 9 were changed. See the MUSE news page.
- July 9: First release of KiDS, the dark matter survey using the OmegaCAM.
- July 8: X-shooter spectra show that some long-duration gamma-ray bursts are driven by magnetars.
- July 7: VISIR is back in operation, as the problem with its astatic support has been fixed.
- June 29: NACO is back in operation, with the camera wheel fixed on S13. However, despite 8 attempts, it was not possible to recover the detector fully: NACO will thus be operated with 3 quadrants only.
- June 25: FLAMES data provides evidence that M87 accreted a ~ 6 x 109 Msun galaxy in the last giga-year.
- June 23: VIRCAM is back in operation. A short power cut today affecting all instruments caused further problems but were quickly solved this time.
- June 19: KMOS is back in operation.
- June 18: UVES is back in operation following the power cut. Problems still affect KMOS. If the detector is fine, NACO could be back in operation on June 19th. VIRCAM should be operational at the earliest on June 21.
- June 17: X-shooter and SINFONI helped to provide the best observational evidence of first generation stars in the Universe.
- June 17: HAWK-I is recovered.
- June 16: FORS2 is back in operation.
- June 15: FLAMES-GIRAFFE, SPHERE, VIMOS, MUSE, and OmegaCAM are back in operation following the power cut.
- June 14: Passively cooled instruments could be back in operation as fast as the rest of the infrastructure allowed, such as X-shooter and SINFONI. However problems still affect UT1, UT3, VST and many systems.
- June 12: A power cut affected the whole observatory: caused by the failure of a back-up generator while the main (gas) turbine was undergoing maintenance. No operation took place on the nights of June 12th and 13th.
- June 10: SPHERE reveals earliest stage of planetary nebula formation.
- May 21: VISIR is warming up in advance of an intervention aimed at fixing a problem with the astatic support found as the cause of the observed drift. It is expected that the instrument would be back in operation on July 7th.
- May 13: FLAMES discovers a new kind of globular star cluster
- May 12: The cause of the drift seen with VISIR has been identified and requires a new intervention to fix the problem.
- April 30: MUSE reveals The Pillars of Creation in 3D
- April 16: Together with Hubble, SINFONI observations show that star formation shuts down in the centres of elliptical galaxies first
- April 15: Has MUSE found the First Signs of Self-interacting Dark Matter?
- April 2: SPHERE enters regular operation. Note however that most DIT/NDITs need to be updated. See the SPHERE News page for details.
- April 2: A drift problem affects observations for most targets with VISIR, limiting objects to bright ones for imaging and excluding spectroscopic observations.
- March 26: SINFONI has provided the Best View Yet of Dusty Cloud Passing Galactic Centre Black Hole
- March 20: NACO: The mechanism that allows to switch objectives suffered a severe failure and CONICA had to be warmed up to be serviced. See more details in the NACO News.
- March 11: VST- OmegaCAM finds A Grand Extravaganza of New Stars
- March 5: Start of the work to refurbush the VLTI Laboratory. MIDI has been decommissioned. Note that the instrument web page has been moved! The VLTI will be offered again in Period 96.
- March 5: CRIRES has mapped the distribution of H2O and HDO on Mars: The Planet that Lost an Ocean’s Worth of Water
- March 04: KMOS: After an intervention to upgrade all arms, the instrument is handed back to Science Operation for recommissioning and testing.
- February 26: MUSE goes beyond Hubble: Looking Deeply into the Universe in 3D
- February 18: First SPHERE Science results: The Strange Case of the Missing Dwarf
- Feburary 9: FLAMES:
- The cleaning of the Giraffe HR grating was again successfully performed at the end of January 2015, resulting in an increase in efficiency estimated to be in the range 10-20% for bluer settings.
- On the afternoon of February 2, 2015, an update of the focussing procedure was implemented in FLAMES/Giraffe. This results in significant increase in both spectral resolution and sensitivity across all settings, although these enhencements vary from setting to setting. See the FLAMES News page.
- February 9: Stellar Partnership Doomed to End in Catastrophe
- February 4: VISTA Stares Right Through the Milky Way
- January 22: KMOS Arm 2 was put in simulation.
- January 17: NACO is resuming operation. The new Aladdin2 (former ISAAC LW) detector installed in CONICA features slightly different noise properties than the Aladdin3. In addition, there is a significant additional thermal background that will slightly reduce the dynamic range and detection limits when long exposure times are used, associated with short wavelength broadband filters. See more details in the NACO News.
- December 4: SPHERE Science Verification has started.
- December 2: KMOS Arm 19 was put into simulation
- November 21: KMOS Arm 8 was put into simulation.
- November 18: KMOS Arm 4 was put into simulation.
- November 15: The FORS2 LADC, which was suffering from a poor coating, was replaced by the one of FORS1 whose coating had been removed. This led to a gain in throughput of 0.1 ( in the blue), 0.07 (in V), 0.05 (in I), and 0.04 (in I) magnitudes, due to a reduction of the scattering. See the FORS2 news page for updates.
- October 6: KMOS Arms 5,6,7,and 16 have been replaced, Electronics of Arm4 were changed. All arms except 11 are back in operation.
- NACO: updates following the move of NACO to UT1 can be seen on the NACO news page.
- August 24: KMOS arm 7 is parked, locked and in simulation
- August 6, 7: KMOS arm 5 is parked, locked and in simulation; arm 15 is recovered and back in operation
- July 17: last night of operation for CRIRES before the start of the upgrade
- July 7: on July 4, an anomalous noise was heard coming from inside the OMEGACAM's cryostat. By precaution, the instrument was out of operation on the nights of July 4 and 5th.
- June 20 - 29: MUSE Science Verification
- June 4: press release describing the first light of SPHERE.
- June 2: speeds of the grating wheel of SINFONI had to be decreased
- May 31: KMOS arm 15 is parked, locked and in simulation
- May 20: KMOS arms 4 and 11 are temporarily out of operations.
- May 20: FLAMES ETC with improved efficiency measures at all settings.
- May 10: start of the SPHERE commissioning.
- April 1: KMOS Arm 11 is back in operation.
- March 5: press release describing the first light of MUSE.
- March 1: Limiting magnitudes for AMBER in the K-band has increased, owing to polarization control.
- February 21: KMOS was successfully recommissioned after an intervention to lower the vibration and to repair the arms that were locked during the last months. Arm 11 is still not available, all other arms work well within the specifications.
- February 7: start of the MUSE commissioning.
- January 7: KMOS arm 11 is parked, locked and in simulation.
- January 7: UVES red arm CCD shutter failure. See the UVES News page.
- December 23: KMOS arm 10 is parked, locked and in simulation
- October 15: KMOS arm 2 is parked, locked and in simulation
- October 1: KMOS is now in operation
- HAWKI on UT4 is now out of service, due to an intervention aimed at cleaning the entrance window. The intervention will be completed by next October 23, 2013.
- March 11: SINFONI: The motors for the Grating and Filter wheels are failing. An intervention is planned for September 2013. In the meantime, users are asked to reduce the number of spectrograph configurations as much as possible. See SINFONI news for details.
- Feburary 27: VISIR is not offered neither in P91 nor in P92, as a better understanding of the detector issues is needed. See the VISIR Upgrade Project page.
- January 17: Start of the 2nd KMOS commissioning.
- December 18: FLAMES-UVES: Fiber 37, plate 2 broke on September 11. Latest attempt to fix it yielded 30% of the original transmission. Therefore the fiber will remain disabled until better results are reached or the fiber is replaced. Follow-up on FLAMES news
- December 4: The first commissioning of KMOS has been successfully completed.
- November 20: NACO intervention is completed. See NACO news for details.
- September 7, 2012: NACO is foreseen to be refurbished from October 1, to mid-November 2012
- September 7, 2012: NACO: Between August 25th and September 5th, issues with the NACO field selector electronics prevented the use of the IR WFS. More details in the NACO news.
- August 16: VISIR: The first commissioning of the upgraded VISIR took place between July 28 and August 10. Results are mixed. The prism provides the expected coverage and resolution. Good progress has been made on the characterisation of the new modes. Unfortunately, the imager does not provide the expected throughput. The instrument will be brought back to the integration hall before the M1 coating (Aug 27 - Sep 6). A search for the cause of the decreased sensitivity will be carried out in addition to planned modifications. It is foreseen that the instrument will not be operational before end of October. More details in the VISIR Upgrade Project page.
- August 16: FORS2: The 'CHIP 1' (or upper chip) of the Red CCD mosaic failed in the morning of August 8, 2012. The Red CCD mosaic was replaced by the Blue one on August 10 to allow for a better set-up diagnostics on the CCD test bench. Several problems were found and solved. The Red CCD mosaic was put back on the instrument on August 16.
- August 16: XSHOOTER: Recently the rate of XSHOOTER ADC failures have increased. There is an ongoing investigation to find the cause for this misbehaviour, but it is unlikely that the situation is back to normal for the next few months. A major intervention on XSHOOTER to fix this problem is under evaluation. Incorrect position of ADCs might lead to slit losses worse than if they are not used. Consequently, the ADCs were temporarily disabled, starting on August 1st, 2012. More details in the XSHOOTER News page.
- June 4: KMOS, the VLT 2nd Generation K-band multi-object spectrograph, will start to be commissioned in November 2012.
- March 21 - May 3: UT4 - Yepun: Because of preparatory work for the installation of the Adaptive Optics Facility, Yepun is not available for operations. Observations with HAWK-I, NACO and SINFONI are therefore not possible.
- February 15, 2012: NACO: the detector of CONICA suffers strong electronic patterns affecting the two upper quadrants. See the NACO News page for details.
- October 15, 2011: OmegaCAM enters in operations.
- October 13, 2011: CRIRES is back in operations following an internvention whose main goal was to replace the entrance slit mechanism of CRIRES. CRIRES has now 2 fixed slits, of widths corresponding to 0.2 and 0.4 arcsec on sky, engraved in a gold coated silicon wafer plate by microlithography and etching.
- July 12, 2011: ISAAC is back: a small piece of metal had managed its way towards the connectors of the Hawaii detector and caused ashort circuit. It could be removed and fortunately the detector was not damaged.
- June 3, 2011: Since May 18, ISAAC can only be operated using the long-wavelength Aladdindetector. The spare Hawaii detector is being tested and should be installed in ISAAC soon. ISAAC should be warmed up on June 6. The instrument should be back in operations by the end of June.
- May 18, 2011: The Hawaii detector used in the ShortWavelength arm of ISAAC has 'lost' one quadrant. The suspected natureof the fault suggests that it should be switched off for safety reason.
- April 28, 2011: The intervention and recommissioning of the NACO visible wavefront sensor were successfull. The Visible WFS is nowback in operations.
- March, 2011: Start of OmegaCAM commissioning.
- March 20: A high-risk earthquake hit Paranal affecting UVES(image slicer alignment), ISAAC (rotator offset changed by 0.1deg).
- February, March. The selector of the visible wavefrontsensor of NACO is unreliable. As of March 6, only the visible 14x14 wavefrontsensor is available, preventing in particular the use of NACO with theLaser Guide Star Facility. An intervention is planned to restore thefullchoice by April 24. The IR WFS is not affected.
- November 22-25: UVES: The turntable for the grating of the UVES blue-arm has been replaced.
- August, September: VIMOS is back on sky, but with problems. See the VIMOS News page.
- May 22 - July 31: VIMOS is taken out of operations forupgrade, in particular, with new red-sensitive detectors and AutomaticFlexure Compensation.
- July 16: CRIRES: A new intervention improved the parallelism of the entrance slit and its alignment with the intermediate slit.
- June: MIDI: The readout windows were shifted by 7 pixels, to recenter the PRISM spectra.
- May 01 until July 10: CRIRES: Problems affecting the coolingmade the instrument unavailable for most of May during which a newcompressor was installed. Unfortunately, following the intervention,the entrance slit lost its parallelism and its alignment with theintermediate slit.
- May: SINFONI: A new STRAP, without central obscuration, was installed. Commissioning hampered by bad weather.
- April 2-3: NACO: The Apodized Phase Plate coronograph was successfully commissioned.
- March 30: CRIRES: The MACAO shutter was replaced.
- Feburary 20: FLAMES-GIRAFFE: Maintenance and repair of CCD shutter.
- Feburary 13-15: ISAAC: Filter wheel 2 was repaired.
- Feburary 7: NACO: Successfull installation and commissioning of a 14x14 Visible WFS for use with the Lase Guide Star Facility.
- October, 2009: XSHOOTER has been fully commissioned and installed at the Cassegrain of UT2; the instrumentofficially started scientific operations from October 1st.
- September, 2009: ISAAC is back in operations after to be re-installed at the Nasmith A of UT3.
- August, 2009: ISAAC has been succesfully installed at the Nasmith A of UT3; the recommissioning on sky is currently finalizing.The Instrument should be back in operations before the end of the month.
- July, 2009: The UVES red-arm has succesfully recommissioned and the instrument is fully back in operations fromJuly 22nd. The UVES red-arm is scheduled to be recomissioned July 10-12 and will be offered onward. ISAAC will be removed from UT1 at the end of this month, and remounted to UT3 early September to help balancing pressure factor across the UTs.
- April, 2009: The two FORSes have ben merged into a single instrument, FORS2,which is still offered at UT1, but with polarimetric capabilities,starting P83. X-SHOOTER has been moved to UT2 for continuation of commissioning and is scheduled to start operations next semester (P84). VISIR has been reinstalled at the UT3 Cassegrain focus and is fully back in operations.
- October, 2008: HAWKI has been fully commissioned and installed at the Nasmith focus of UT4; the instrument officially startedscientific operations from October 1st.
- August 1st, 2007: HAWKI: First Light of the instrument mounted at UT4.
- April 1st, 2007: AMBER: start of Science Operation with 3 ATs. Fringe tracking with FINITO is availabe to visitor mode PIs
- April 1st, 2007: CRIRES: start of Science Operation
- April 7, 2007: FORS1 CCD upgrade completed
- March 25, 2007: First light for MAD, the multi-conjugate adaptive optics demonstrator
- March 13, 2007: SINFONI: first commissioning with LGS completed
- March 8, 2007: Installation of new FLAMES FACB guiding TCCDs completed
- January 2007: NACO: first commissioning with LGS completed
- December 18, 2006: LGSF successfully commissioned
- December 5, 2006: Installation of new UVES slit viewers TCCDs completed
- October 31, 2006: DAZLE visitor instrument succesfully commissioned on UT3
- October 29, 2006: Launch Telescope of LGSF successfully re-installed
- June 2006: commissioning of CRIRES at UT1 started
- April 2006: CRIRES AO system (MACAO) installed and commissioned at UT1
- January 28, 2006: first light of the Laser Guide Star Facility LGSF at UT4
- October 01, 2005: AMBER: start of Science Operation with 3 UTs
- October 01, 2005: MIDI: start of Science Operation with 2 ATs
- May 04, 2005: first light for the ULTRACAM visitor instrument at the Nasmyth A focus of Melipal (UT3)
- March 24, 2005: SINFONI early start of Science Operation
- March 20, 2005: VISIR early start of Science Operation
- March 01, 2005: New and uniform instrument webpages released
- February 2005: VINCI and MIDI: first fringes with two ATs
- January 2005: SINFONI Science Verification Data available for the August, October, and December 2004 runs.
- January 2005: VISIR Science Verification Data available for the runs between September and December 2004.
- October 20, 2004: UVES: detector in the blue arm exchanged
- October 1, 2004: Rapid Response Mode (RRM) available for FORS1, FORS2, ISAAC, and UVES
- July 9, 2004: SINFONI: first light for SPIFFI with the AO module
- July 2004: MIDI Science Demonstration Data Release
- June 1, 2004: SINFONI: first light for AO module on Yepun (UT4, Cassegrain focus)
- June 1, 2004: AMBER: first fringes with three UTs
- June 2004: NACO: New Conica detector (Aladdin3) available
- June 2004: FORS1 moved from Antu (UT1) to Kueyen (UT2)
- May 2004: FORS2 moved from Yepun (UT4) to Antu (UT1)
- May 28, 2004: AMBER: first fringes with two UTs
- April 30, 2004: VISIR: first light on Melipal (UT3, Cassegrain focus)
- April 1, 2004: Rapid Response Mode (RRM) available for UVES
- April 1, 2004: MIDI: start of Science Operations
- March 20, 2004: AMBER: first fringes with two siderostats
On December 16 at 05:10 LT another earthquake with magnitude 6.8 ( Richter) affected the north of Chile. In Paranal it was registered with a magnitude of 5.2 and a peak ground acceleration of 0.2m/s2. Neither injuries to the people nor damages to the installations have been registered.
On November 14 at 12:41 LT an earthquake with magnitude 7.7 (Richter) with epicenter north of Antofagasta affected the north of Chile. In Paranal it was registered with a magnitude of 5.7, a peak ground acceleration of 0.6m/s2 and for a duration of approx 2 minutes. This has been the strongest earthquake since 1997, the first "high risk" earthquake since the start of operation of the VLT, and the first strong earthquake since the opening of the residencia. Neither injuries to the people nor damages to the installations have been registered. After extended checkout procedures, the regular night time operations have been resumed at the sunset. All the systems have been operated without restrictions, including all four UTs (with their instruments and the LGS) and the ATs with VLTI.
Several other minor episodes have been registered in the following 24 hours without any impact on the normal VLT activities.